How many grams of lidocaine total are in 200 mL of 2% lidocaine solution added to 300 mL of 3% lidocaine solution
Answer by stanbon(75887)

How many grams of lidocaine total are in 200 mL of 2% lidocaine
solution added to 300 mL of 3% lidocaine solution
----
lido = 0.02*200 + 0.03*300 = 4 + 9 = 13 mL of lido
------------
Cheers,
Stan H.
============